# 3D Gaussian Parametric Head Model

### v1

@@ Abstract
Creating high-fidelity 3D human head avatars is crucial for applications in VR/AR, telepresence, digital human interfaces, and film production. Recent advances have leveraged morphable face models to generate animated head avatars from easily accessible data, representing varying identities and expressions within a low-dimensional parametric space. However, existing methods often struggle with modeling complex appearance details, e.g., hairstyles and accessories, and suffer from low rendering quality and efficiency. This paper introduces a novel approach, 3D Gaussian Parametric Head Model, which employs 3D Gaussians to accurately represent the complexities of the human head, allowing precise control over both identity and expression. Additionally, it enables seamless face portrait interpolation and the reconstruction of detailed head avatars from a single image. Unlike previous methods, the Gaussian model can handle intricate details, enabling realistic representations of varying appearances and complex expressions. Furthermore, this paper presents a well-designed training framework to ensure smooth convergence, providing a guarantee for learning the rich content. Our method achieves high-quality, photo-realistic rendering with real-time efficiency, making it a valuable contribution to the field of parametric head models.
Keywords:
3D Gaussian Head Avatar Parametric Model
@@ 1 Introduction
Creating high-fidelity 3D human head avatars holds significant importance across various fields, including VR/AR, telepresence, digital human interfaces, and film production. The automatic generation of such avatars has been a focal point in computer vision research for many years. Recent methods [61, 56, 55, 13, 65, 62, 63, 17, 12, 38] can create an animated head avatar through conveniently collected data such as monocular video data or even a picture [26, 22]. Serving as the most fundamental tool in these methods, the 3D morphable models (3DMM) [14, 25], which represent varying identities and expressions within a low-dimensional space, have been proven to be a highly successful avenue in addressing this challenging problem.
Since the traditional parametric 3DMMs are typically limited by the topology of the underlying template mesh and only focus on the face part, some works [59, 28, 15, 16] propose to use implicit Signed Distance Field (SDF) as the geometric representation to model the entire head. Despite their flexibility, these methods fall short in recovering high-frequency geometric and texture details like hairstyles, glasses or accessories. On the other end of the spectrum, Neural Radiance Field (NeRF) [34] based methods [64, 19] learn parametric head models by directly synthesizing images, thus eliminating the need of geometry modeling. However, NeRF is built upon volumetric rendering, which involves sampling and integrating points distributed throughout space. Therefore, NeRF-based methods typically suffer from low rendering efficiency and have to trade it off with rendering resolution, thereby greatly reducing rendering quality. Moreover, skipping geometric reconstruction would probably lead to poor 3D consistency.
More recently, 3D Gaussian Splatting (3DGS) [21], which uses explicit Gaussian ellipsoids to represent 3D scenes, has attracted significant attention from the research community. Experiments have verified the superior quality of the rendered results and excellent rendering efficiency compared to previous NeRF-based or surface-based methods even on dynamic scenes [48, 57, 32, 58]. Motivated by this progress, we propose a novel 3D Gaussian Parametric Head Model, which, for the first time, marries the power of 3DGS with the challenging task of parametric head modeling. Our 3D gaussian parametric head model decouples the control signals of the head into the latent spaces of identity and expression, as is also done in SDF-based face model NPHM [15]. These latent spaces are then mapped to the offsets of the Gaussian positions, which effectively represent the variance of shape and appearance of different identities and expressions. Benefiting from the differentiability of Gaussian splatting, our model can be learned from multi-view video data corpus in an end-to-end manner, without relying on geometry supervision.
@@
After training on large corpus of multi-view head videos, our parametric Gaussian head model can generate photorealistic images that accurately depict the diverse range of facial appearances, naturally handling complex and exaggerated expressions, while also enabling real-time rendering. Additionally, our method supports single-image fitting and surpasses previous techniques in both reconstruction accuracy and identity consistency. Furthermore, the model resulting from our fitting process allows for the control of various expressions while maintaining naturalness and consistent identity even under exaggerated expressions. The contributions of our method can be summarized as:
  • We propose 3D Gaussian Parametric Head Model, a novel parametric head model which utilizes 3D Gaussians as the representation and enables photo-realistic rendering quality and real-time rendering speed.
@@
  • Our 3D Gaussian Parametric Head Model enables the generation of a detailed, high-quality face avatar from a single given image, as well as performing expression and identity editing upon it.
@@ 2 Related Work
3D GAN based Head Models. 3D Generative Adversarial Networks (GANs) have revolutionized the field of computer vision, particularly in the domain of human head and face modeling, enabling the generation of face avatars from input images. Traditional methods often require labor-intensive manual work or rely on multi-view images to create 3D models. 3D GANs as a more automated and data-driven approach, which are just trained on single-view 2D images but generate detailed and realistic 3D models of human head [8, 7, 18, 35, 11, 52, 9]. Panohead [1] additionally introduces images of hairstyles on the back of characters and trains a full-head generative model. Based on the previous methods, IDE-3D [42] proposes to use semantic map to edit the 3D head model. Next3D [43] and AniFaceGAN [50] extend to use the FLAME model [25] to condition the generated head model, so that the expression and pose of the generated head model can be controlled. AniPortraitGAN [51] further replaces FLAME model with SMPLX model [36] to generate upper body avatars, thus the shoulders and the neck can also be controlled. These 3D GAN-based models primarily leverage the coarse FLAME model for expression control, often leading to a loss of expression details in the generated faces. In contrast, our method directly learns the expression distribution from the dataset, capturing more facial appearance details.
@@
3D Gaussians. Recently, 3D Gaussian splatting [21] has shown superior performance compared to NeRF, excelling in both novel view synthesis quality and rendering speed. Several methods have expanded Gaussian representation to dynamic scene reconstruction [48, 57, 32, 58]. For human body avatar modeling, recent approaches [27, 20] propose training a 3D Gaussian avatar animated by SMPL [31] or a skeleton from multi-view videos, surpassing previous methods in rendering quality and efficiency. In the realm of human head avatar modeling, recent techniques [54, 37, 39, 45, 23, 37, 10, 33, 53, 40] also utilize 3D Gaussians to create high-fidelity and efficient head avatars. These approaches center on the creation of a high-fidelity person-specific avatar using data of a single person. In contrast, our method focuses on a versatile prior model that can accommodate varying appearances. Once trained, our model is also capable of person-specific avatar reconstruction by fitting to the input image data provided.
@@ 3 Method
Figure 2: The pipeline of our method. Our training strategy can be divided into a Guiding Geometry Model for initialization, and a final 3D Gaussian Parametric Head Model. Deformations of each model are further decoupled into identity-related and expression-related deformations. Rendering involves using DMTet to transform the initial model into a mesh and 3D Gaussian Splatting for the Gaussian model. Features from both models are finally upsampled to high-resolution portrait images through a convolutional network $\boldsymbol{\Psi}$. During inference, our output exclusively comes from the Gaussian model.
In this section, we present the 3D Gaussian Parametric Head Model. In contrast to previous mesh-based or NeRF-based models, initializing and training Gaussian-based models pose distinct challenges. This section introduces the dataset and preprocessing, the carefully designed guiding geometry model, the Gaussian Parametric Head Model, and outlines their respective training processes. Additionally, we will also provide the training details and demonstrate how to utilize our method when given a single input image.
@@ 3.1 Data Preprocessing
We used three datasets for our model training, including a multi-view video dataset NeRSemble [24], and two 3D scans datasets NPHM [15] and FaceVerse [47]. We do not use the 3D geometry of the scans directly, but render them into multi-view images and use only the images form the 3 datasets as supervision. In order to better utilize these three different datasets, we need to do preprocessing. First, we resize the images to 512 resolution and adjust the camera parameters. Then, we use BackgroundMattingV2 [29] to extract the foreground characters in the NeRSemble dataset and record the masks. This step is not required for the two synthetic datasets. Next, we use face alignment [4] to detect 2D landmarks in all the images. Through these 2D landmarks, we fit a Basel Face Model (BFM) [14] for each expression of each identity, and record the head pose and 3D landmarks of the BFM. We will use the above processed camera parameters, images, masks, head pose of BFM and 3D landmarks of BFM to train our model.
@@ 3.2 Model Representation
The representation of Gaussian distribution poses challenges due to its unordered and unstructured nature, leading to difficulties in the continuous spread of gradients to neighboring points in space during backpropagation. This often results in convergence failure when Gaussians are randomly initialized. On the other hand, surface-based representations such as mesh are just suitable for rough geometry learning. A direct idea is to utilize an existing 3DMM, such as FLAME [25], as the initial position for the points in 3D Gaussian splatting [21]. However, this coarse initialization still fails to converge the positions of 3D points to the correct locations, as shown in Fig. 4. The network tends to alter the shape of the ellipsoid to achieve a suitable fitting result, leading to inaccurate geometry of the point cloud and blurriness in the rendered image.
@@
Guiding Geometry Model. The guiding geometry model receives an identity code $\boldsymbol{z^{id}}$ and an expression code $\boldsymbol{z^{exp}}$ as input, producing a mesh with vertices $V$, faces $F$, and per-vertex color $C$ that aligns with the specified identity and expression. To achieve this, we use an MLP denoted as $\boldsymbol{f_{mean}}(\cdot)$ to implicitly model the SDF, which represents the mean geometry:
@@
where $s$ denotes the SDF value, $\gamma$ denotes the feature from the last layer and $x$ denotes the input position. Then, we convert the implicit SDF through Deep Marching Tetrahedra (DMTet) [41] into an explicit mesh with vertex positions $V_{0}$, per-vertex feature $\Gamma$ and faces $F$. Next, we need to transform the mean shape into a neutral-expression shape on condition of the input identity code $\boldsymbol{z^{id}}$. To inject identity information into the vertices of the mesh, we first use an injection MLP $\boldsymbol{f_{inj}}(\cdot)$, which takes the identity code $\boldsymbol{z^{id}}$ and the per-vertex feature $\Gamma$ as input and produces the identity-conditioned per-vertex feature vectors $H=\boldsymbol{f_{inj}}(\boldsymbol{z^{id}},\Gamma)$. Subsequently, utilizing a tiny MLP $\boldsymbol{f_{id}}(\cdot)$, we predict the displacement $\delta V_{id}$ for each vertex. This displacement is used to transform the mean shape into the neutral-expression shape conditioned on the id code $\boldsymbol{z^{id}}$.
@@
After completing deformations related to identity, the next step is to capture the deformation induced by facial expressions. We introduce another tiny MLP $\boldsymbol{f_{exp}}(\cdot)$. This MLP takes the feature vectors $H$ obtained in the previous step and the expression code $\boldsymbol{z^{exp}}$ as input, and the output is the displacement $\delta V_{exp}$ for each vertex. Using this displacement, we update the vertex positions to $V_{can}$. Additionally, we feed the same feature vectors $H$ and expression code $\boldsymbol{z^{exp}}$ to a color MLP, $\boldsymbol{f_{col}}(\cdot)$, to predict the 32-channel color $C$ for each vertex. The vertex positions to $V_{can}$ and 32-channel color $C$ can be described as:
$V_{can}=V_{0}+\boldsymbol{f_{id}}(H)+\boldsymbol{f_{exp}}(H,\boldsymbol{z^{exp}}),\ C=\boldsymbol{f_{col}}(H,\boldsymbol{z^{exp}}).$ (2)
[…]
@@
3D Gaussian Parametric Head Model. The Gaussian model also takes an identity code $\boldsymbol{z^{id}}$ and an expression code $\boldsymbol{z^{exp}}$ as input, producing the positions $X$, color $C$, scale $S$, rotation $Q$ and opacity $A$ of the 3D Gaussians. Similar to the guiding geometry model, we initially maintain an overall mean point cloud, with the mean positions $\boldsymbol{X_{0}}$. However, we no longer generate the per-vertex feature $\Gamma$ through $\boldsymbol{f_{mean}}(x)$. Instead, we directly generate it at once and bind it to the Gaussian points as optimizable variables $\boldsymbol{\Gamma_{0}}$. This is possible since the number of Gaussian points is fixed at this stage. Then we need to transform the mean point cloud into a neutral-expression point cloud, conditioned by the id code $\boldsymbol{z^{id}}$. To achieve this, we utilize the same injection MLP $\boldsymbol{f_{inj}}(\cdot)$ and identity deformation MLP $\boldsymbol{f_{id}}(\cdot)$ defined in the guiding geometry model, which can generate feature vectors $H=\boldsymbol{f_{inj}}(\boldsymbol{z^{id}},\boldsymbol{\Gamma_{0}})$ that encode identity information for each point and predict the identity-related displacement of each point. Then, we also need to predict the expression code $\boldsymbol{z^{exp}}$-conditioned displacement. The resulting positions $X_{can}$ and the 32-channel color $C$ of each point, similar to the approach presented in the guiding geometry model, can be described as:
$X_{can}=\boldsymbol{X_{0}}+\boldsymbol{f_{id}}(H).+\boldsymbol{f_{exp}}(H,\boldsymbol{z^{exp}}),\ C=\boldsymbol{f_{col}}(H,\boldsymbol{z^{exp}}).$ (3)

### v2

# GPHM: Gaussian Parametric Head Model for Monocular Head Avatar Reconstruction

@@ Abstract
Creating high-fidelity 3D human head avatars is crucial for applications in VR/AR, digital human, and film production. Recent advances have leveraged morphable face models to generate animated head avatars from easily accessible data, representing varying identities and expressions within a low-dimensional parametric space. However, existing methods often struggle with modeling complex appearance details, e.g., hairstyles, and suffer from low rendering quality and efficiency. In this paper we introduce a novel approach, 3D Gaussian Parametric Head Model, which employs 3D Gaussians to accurately represent the complexities of the human head, allowing precise control over both identity and expression. The Gaussian model can handle intricate details, enabling realistic representations of varying appearances and complex expressions. Furthermore, we presents a well-designed training framework to ensure smooth convergence, providing a robust guarantee for learning the rich content. Our method achieves high-quality, photo-realistic rendering with real-time efficiency, making it a valuable contribution to the field of parametric head models. Finally, we apply the 3D Gaussian Parametric Head Model to monocular video or few-shot head avatar reconstruction tasks, which enables instant reconstruction of high-quality 3D head avatars even when input data is extremely limited, surpassing previous methods in terms of reconstruction quality and training speed. Project page: https://yuelangx.github.io/gphmv2/.
Index Terms:
Gaussian Splatting, Parametric Model, Head Avatar
@@ 1 Introduction
Creating high-fidelity 3D human head avatars holds significant importance across various fields, including VR/AR, telepresence, digital human interfaces, and film production. The automatic generation of such avatars has been a focal point in computer vision research for many years. Recent methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] can create an animated head avatar through conveniently collected data such as a monocular video data or even a picture [11, 12]. Serving as the most fundamental tool in these methods, the 3D morphable models (3DMM) [13, 14], which represent varying identities and expressions within a low-dimensional space, have been proven to be a highly successful avenue in addressing this challenging problem.
Since the traditional parametric 3DMMs are typically limited by the topology of the underlying template mesh and only focus on the face part, some works [15, 16, 17, 18] propose to use implicit Signed Distance Field (SDF) as the geometric representation to model the entire head. Despite their flexibility, these methods fall short in recovering high-frequency geometric and/or texture details like complex hairstyles, glasses or accessories. On the other end of the spectrum, Neural Radiance Field (NeRF) [19] based methods [20, 21] learn parametric head models by directly synthesizing photo-realistic images, thus eliminating the need of geometry modeling. However, NeRF is built upon volumetric rendering, which involves sampling and integrating points distributed throughout space. Therefore, NeRF-based methods typically suffer from low rendering efficiency and have to trade it off with rendering resolution, thereby greatly reducing rendering quality. Moreover, skipping geometric reconstruction would probably lead to poor 3D consistency.
More recently, 3D Gaussian Splatting (3DGS) [22], which uses explicit Gaussian ellipsoids to represent 3D scenes, has attracted significant attention from the research community. Experiments have verified the superior quality of the rendered results and excellent rendering efficiency compared to previous NeRF-based or surface-based methods even on dynamic scenes [23, 24, 25, 26]. Motivated by this progress, we propose a novel 3D Gaussian Parametric Head Model for head avatar modeling, which, for the first time, marries the power of 3DGS with the challenging task of parametric head modeling. Our 3D gaussian parametric head model decouples the control signals of the head into the latent spaces of identity and expression, as is also done in SDF-based face model NPHM [17]. These latent spaces are then mapped to the offsets of the Gaussian positions, which effectively represent the variance of shape and appearance of different identities and expressions. Benefiting from the differentiability of Gaussian splatting, our model can be learned from multi-view video data corpus in an end-to-end manner, without relying on geometry supervision, achieving high quality monocular head avatar reconstruction results.
@@
After training on large corpus of multi-view head videos, our parametric Gaussian head model can generate photorealistic images that accurately depict the diverse range of facial appearances, naturally handling complex and exaggerated expressions, while also enabling real-time rendering. Additionally, our method supports single-image fitting and surpasses previous techniques in both reconstruction accuracy and identity consistency. Furthermore, the model resulting from our fitting process allows for the control of various expressions while maintaining naturalness and consistent identity even under exaggerated expressions.
A preliminary version of this work has been published in ECCV 2024 [27], in which we propose a novel 3D Gaussian Parametric Head Model (GPHM) enabling photo-realistic representation of human heads and high-quality face avatar from a single image. However, the preliminary work [27] mainly focuses on representing a parametric head model, lacking ease of use and robustness for downstream tasks like head reconstruction from input images. In the current version, we present GPHMv2, a head avatar reconstruction framework, which supports instant and robust head avatar reconstruction from monocular video or even few-shot image inputs. The proposed head reconstruction pipeline surpasses previous NeRF-based [3, 4, 5, 2] or 3DGS-based [28, 29] head reconstruction methods in reconstruction quality and training speed. Moreover, previous methods heavily rely on the 3DMM models [14, 13], suffering from the coupling of expression and shape, and perform poorly in cross-identity reenactment (see Sec 4). And due to lacking sufficient prior information, these methods can hardly support few-shot or one-shot head reconstruction like our method.
We extend the preliminary version [27] as follows. Firstly, we extend our network structure and adjust data preprocessing for a more expressive and generalizable head model. Specifically, to better disentangle the expression and head motion of the avatar and capture more detailed expression information, we introduce a facial expression encoder and a non-face motion encoder to extract latent expressions and latent motions from images (see Section 3.23.3). These components are jointly trained with the other network components to enable end-to-end avatar animation via image reconstruction results. However, training the model in an end-to-end self-reconstruction manner and directly using the input image as the expression condition inevitably leads to the leakage of identity-related appearance information into the latent expression codes. To address this, we utilize LivePortrait [30] to synthesize a large number of images with different identities but the same expression as the additional expression condition during training, effectively eliminating this issue (see Section 3.1). Therefore, the model can more accurately isolate and capture expression features independently of identity, leading to more precise and versatile avatar animations.
Secondly, we enhance the functionality of the preliminary model by designing a few-shot 3D head avatar reconstruction framework. Leveraging the pre-trained extended GPHMv2 model, a high-quality 3D head avatar can be rapidly reconstructed using only a small amount of monocular data. Specifically, we optimize our model for a single identity-specific avatar in a two-stage process, from coarse to fine. In the first stage, we focus solely on optimizing the identity code to provide a rough initialization. In the second stage, we refine the model by optimizing the neutral Gaussian attributes and motion-related networks to capture finer details. Also in this stage, we introduce a tiny network to accurately model the expression-related dynamic changes in color and Gaussian attributes. Finally, the well-finetuned head avatar can be driven by any video fed into the encoders.
The contributions of our method can be summarized as:
  • We propose 3D Gaussian Parametric Head Model, a novel parametric head model which utilizes 3D Gaussians as the representation, enabling photo-realistic rendering quality and real-time rendering speed.
@@
  • We extend our preliminary proposed 3D Gaussian Parametric Head Model to the more expressive and generalizable GPHMv2, which enables instant head avatar reconstruction from monocular video or even few-shot images, achieving state-of-the-art quality and training time.
@@ 2 Related Work
3D GAN based Head Models. 3D Generative Adversarial Networks (GANs) have revolutionized the field of computer vision, particularly in the domain of human head and face modeling, enabling the generation of face avatars from input images. Traditional methods often require labor-intensive manual work or rely on multi-view images to create 3D models. 3D GANs as a more automated and data-driven approach, which are just trained on single-view 2D images but generate detailed and realistic 3D models of human head [39, 40, 41, 42, 43, 44]. Panohead [45] additionally introduces images of hairstyles on the back of characters and trains a full-head generative model. Based on the previous methods, IDE-3D [46] proposes to use semantic map to edit the 3D head model. Next3D [47] and AniFaceGAN [48] extend to uses the FLAME model [14] to condition the generated head model, so that the expression and pose of the generated head model can be controlled. AniPortraitGAN [49] further replaces FLAME model with SMPLX model [50] to generate upper body avatars, thus the shoulders and the neck can also be controlled. These 3D GAN-based models primarily leverage the coarse FLAME model for expression control, often leading to a loss of expression details in the generated faces. In contrast, our method directly learns the expression distribution from the dataset, capturing more facial appearance details.
@@
3D Gaussians-based Head Models. Recently, 3D Gaussian splatting [22] has shown superior performance compared to NeRF, excelling in both novel view synthesis quality and rendering speed. Several methods have expanded Gaussian representation to dynamic scene reconstruction [23, 24, 25, 26]. For human body avatar modeling, recent approaches [51, 52] propose training a 3D Gaussian avatar animated by SMPL [53] or a skeleton from multi-view videos, surpassing previous methods in rendering quality and efficiency. In the realm of human head avatar modeling, recent techniques [54, 55, 56, 57] also utilize 3D Gaussians to create high-fidelity and efficient head avatars. These approaches centers on the creation of a high-fidelity person-specific avatar using data of a single person. In contrast, our method focus on a versatile prior model that can accommodate varying appearances. Once trained, our model is also capable of person-specific avatar reconstruction by fitting to the input image data provided.
Monocular 3D Head Avatar Reconstruction. 3D head avatars reconstruction from monocular videos is also a popular yet challenging research topic. Early methods [58, 59, 60, 61, 62, 63] optimize a morphable mesh to fit the training video. Recent methods [8, 64] leverage neural networks to learn non-rigid deformation upon 3DMM face templates [14, 13], thus can recover more dynamic details. However, such methods are not flexible enough to handle complex topologies. IMavatar [6] proposes to learn head avatars with implicit SDF-based geometry [65, 66], thus getting rid of the topology limitation of the mesh templates. PointAvatar [7] combines the explicit point cloud with the implicit representation to improve the quality of the rendered images. As NeRF [19] demonstrates its ability to synthesize high-fidelity novel view images, several methods [67, 68, 9, 69, 70, 2, 10] attempt to exploit such representation for neural head modeling. Furthermore, the voxel-based data structure is introduced for training acceleration [4, 3, 5]. However, these representations usually suffer from the loss of high-frequency details. To overcome, recent methods [28, 29, 71] introduce 3D Gaussian representation [22] to model the head avatars, thereby improving the reconstruction quality while reducing the training time requirement. However, despite these advancements, there is still potential for further enhancement in reconstruction quality and training/inference speed.
@@ 3 Method
Figure 2: The overview of our GPHM model. Our training strategy can be divided into a Guiding Geometry Model for initialization, and a final 3D Gaussian Parametric Head Model. Deformations of each model are further decoupled into identity-related, expression-related and non-face deformations. For the expression condition images, we input crop groundtruth face image or synthesized images via LivePortrait [30]. For the non-face motion condition, we input groundtruth images with the face area masked. The renderer involves a convolutional refine network $\boldsymbol{\Psi}$, which finally transfers the feature maps from mesh/Gaussian renderer to fine portrait images. During inference, our output exclusively comes from the Gaussian model.
In this section, we present the 3D Gaussian Parametric Head Model. In contrast to previous mesh-based or NeRF-based models, initializing and training Gaussian-based models pose distinct challenges. This section introduces the dataset and preprocessing, the carefully designed guiding geometry model, the Gaussian Parametric Head Model, and outlines their respective training processes. Additionally, we will provide the training details and demonstrate how to utilize our model for head avatar reconstruction when given an input monocular video.
Figure 3: We generate additional expression condition images via LivePortrait [30] for training the appearance decoupled expression encoder.
@@ 3.1 Data Preprocessing
We used 4 datasets for our model training, including a multi-view video dataset NeRSemble [72], a large-scale monocular video dataset VFHQ [73], two 3D scans datasets NPHM [17] and FaceVerse [33]. We do not use the 3D geometry of the scans directly, but render them into multi-view images and use only the images from the 4 datasets as supervision. To better utilize these 4 different datasets, preprocessing is necessary. First, we resize the images to 512 resolution and adjust the camera parameters. Note that for the monocular videos in VFHQ dataset, we assign a global default value to the camera parameters.
Then, we use BackgroundMattingV2 [74] to extract the foreground characters in the NeRSemble dataset and record the masks. For the VFHQ dataset, we use RobustVideoMatting [75] to segment foreground and masks. This step is not required for the two synthetic datasets. Next, we use face alignment [76] to detect 2D landmarks in all the images. Through these 2D landmarks, we fit a Basel Face Model (BFM) [13] for each expression of each identity, and record the head pose and 3D landmarks of the BFM. For all the images in which the facial area is visible, we extract the face region images through the 2D landmarks and the non-face images by masking the face region.
Finally, we use LivePortrait [30] to synthesize additional expression condition images. Specifically, for each frontal face image in the training set, we randomly sample another from VFHQ. The original image provides the expression as the driving image, and the sampled image provides the appearance as the source image, creating a new image with the same expression but a random identity. All synthesized images are used as additional expression condition images to train the expression encoder for better generalization as described in Section 3.2.
We will use the processed camera parameters, images, masks, head pose, 3D landmarks, face images, non-face images, and synthesized expression condition images mentioned above to train our parametric head model.
@@ 3.2 Model Representation
The representation of 3D Gaussians poses challenges due to its unordered and unstructured nature, leading to difficulties in the continuous spread of gradients to neighboring points in space during backpropagation. This often results in convergence failure when Gaussians are randomly initialized. On the other hand, surface-based representations such as mesh are just suitable for rough geometry learning. A direct idea is to utilize an existing 3DMM, such as FLAME [14], as the initial position for the points in 3D Gaussian splatting [22]. However, this coarse initialization still fails to converge the positions of 3D points to the correct locations, as shown in Fig. 6. The network tends to alter the shape of the ellipsoid to achieve a suitable fitting result, leading to inaccurate geometry of the point cloud and blurriness in the rendered image.
@@
In addition, setting a separate expression code for each frame of data to model dynamic motion like the preliminary version [27] will cause facial expressions to be coupled with body parts. And as more data is added, it becomes more difficult to optimize a large number of discrete latent codes.
Therefore, we first use two separate latent codes: facial expression codes and non-face motion codes to control facial expressions and non-facial movements respectively. Furthermore, we introduce two additional motion encoders accordingly to avoid directly optimizing those latent codes. A face encoder extracts facial expression codes from face images, while a non-face encoder extracts non-face motion codes from non-face images. Moreover, the 2 encoders can be directly used to extract motion from driving images via a single forward pass in subsequent tasks.
Guiding Geometry Model. The guiding geometry model receives an identity code $\boldsymbol{z^{id}}$, a face image $I_{face}$ for expression condition and a non-face image $I_{nonface}$ for controlling non-face area as input, producing a mesh with vertices $V$, faces $F$, and per-vertex color $C$ that aligns with the specified identity and expression. To achieve this, we use an MLP denoted as $\boldsymbol{f_{mean}}(\cdot)$ to implicitly model the SDF, which represents the mean geometry:
@@
where $s$ denotes the SDF value, $\gamma$ denotes the feature from the last layer and $x$ denotes the input position. Then, we convert the implicit SDF through Deep Marching Tetrahedra (DMTet) [77] into an explicit mesh with vertices positions $V_{0}$, per-vertex feature $\Gamma$ and faces $F$. Next, we need to transform the mean shape into a neutral-expression shape on condition of the input identity code $\boldsymbol{z^{id}}$. To inject identity information into the vertices of the mesh, we first use an injection MLP $\boldsymbol{f_{inj}}(\cdot)$, which takes the identity code $\boldsymbol{z^{id}}$ and the per-vertex feature $\Gamma$ as input and produces the identity-conditioned per-vertex feature vectors $H=\boldsymbol{f_{inj}}(\boldsymbol{z^{id}},\Gamma)$. Subsequently, utilizing a tiny MLP $\boldsymbol{f_{id}}(\cdot)$, we predict the displacement $\delta V_{id}$ for each vertex. This displacement is used to transform the mean shape into the neutral-expression shape conditioned on the id code $\boldsymbol{z^{id}}$.
$\delta V_{id}=\boldsymbol{f_{id}}(H).$ (2)
@@
After completing deformations related to identity, the next step is to capture the deformation induced by facial expressions and head pose. Here, the same as GHA [54], we define the human face area as the canonical reference system. In addition to facial expression changes, we need to consider the movement of non-face areas such as the neck and body relative to the head while the face is rigidly transformed. Specifically, we introduce 2 motion encoders $\boldsymbol{E}_{exp}(\cdot)$ for face area, $\boldsymbol{E}_{nf}(\cdot)$ for non-face area, and 2 tiny MLPs $\boldsymbol{f_{exp}}(\cdot)$ for face area and $\boldsymbol{f_{nf}}(\cdot)$ for non-face area. The facial expression encoder takes the condition face image $I_{face}$ as input and predict the facial expression code $z^{exp}$. The non-face motion encoder takes the condition non-face image $I_{nonface}$ as input and predict the non-face motion code $z^{nf}$. Note, during training, the face image can be from any one view in the current frame or the synthesized expression condition images. Then, $\boldsymbol{f_{exp}}(\cdot)$ takes the feature vectors $H$ obtained in the previous step and the expression code $z^{exp}$ from the expression encoder as input, and outputs the displacement $\delta V_{exp}=\boldsymbol{f_{exp}}(H,z^{exp})$ for each vertex. $\boldsymbol{f_{nf}}(\cdot)$ takes the feature vectors $H$ and a non-face motion code $z^{nf}$ as input, and outputs the displacement $\delta V_{nf}=\boldsymbol{f_{nf}}(H,z^{nf})$ for each vertex. Using this displacement, we update the vertex positions to $V_{can}$. Additionally, we feed the same feature vectors $H$ to a color MLP $\boldsymbol{f_{col}}(\cdot)$, to predict the 32-channel color $C$ for each vertex. The vertex positions to $V_{can}$ and 32-channel color $C$ can be described as:
$V_{can}$ $=V_{0}+\delta V_{id}+\lambda_{exp}(V_{0})\delta V_{exp}+\lambda_{nf}(V_{0})\delta V_{nf},$ (3)
$C$ $=\boldsymbol{f_{col}}(H).$ (4)
$\lambda_{exp}(\cdot)$ and $\lambda_{nf}(\cdot)$ respectively indicate whether the vertices belong to the face area, affected by the facial expression code, or belong to the non-face area, affected by the non-face motion code.
Here, we assume that the vertices closer to the 3D landmarks are more affected by the expression code and less affected by the non-face motion code, while the opposite is true for the vertices far away. Specifically, The 3D landmarks $\boldsymbol{P}_{0}$ of the canonical model are first estimated through the 3DMM model in the data preprocessing 3.1 and then optimized later. We calculate the above weight $\lambda_{exp}(\cdot)$ and $\lambda_{nf}(\cdot)$ as follows:
$\begin{split}\lambda_{exp}(x)=\left\{\begin{array}[]{ll}1,&dist(x,\boldsymbol{P}_{0})<t_{1}\\ \frac{t_{2}-dist(x,\boldsymbol{P}_{0})}{t_{2}-t_{1}},&dist(x,\boldsymbol{P}_{0})\in[t_{1},t_{2}]\\ 0,&dist(x,\boldsymbol{P}_{0})>t_{2}\end{array}\right.\end{split}$
with $\lambda_{nf}(x)=1-\lambda_{exp}(x)$. And $x\in V_{0}$ denotes the position of one vertex. $dist(x,\boldsymbol{P}_{0})$ denotes the minimum distance from the point $x$ to the 3D landmarks $\boldsymbol{P}_{0}$. $t_{1}=0.1$ and $t_{2}=0.12$ are predefined hyperparameters when the length of the head is set to approximately $1$.
In practice, we find it difficult and unnecessary to learn expression-related color changes in a generalizable head model, as feeding the expression code to the color MLP may lead to the coupling of appearance and expression. Therefore, we do not consider expression-related color changes during the head model training stage, but model such changes in the downstream reconstruction task (see Section 3.5).
Finally, we utilize the estimated head pose parameters $R$ and $T$ obtained during data preprocessing to transform the mesh from the canonical space to the world space $V=R\cdot V_{can}+T$. After generating the final vertex positions, colors, and faces $\{V,C,F\}$ of the mesh, we render the mesh into a 512-resolution 32-channel feature map $I_{F}$ and a mask $M$ through differentiable rasterization with a given the camera pose. Subsequently, the feature map is interpreted as a 512-resolution fine RGB $I_{fine}$ image through a lightweight convolutional refine network $\boldsymbol{\Psi}(\cdot)$, as shown in Fig. 2.
@@
3D Gaussian Parametric Head Model. The Gaussian model also takes an identity code $\boldsymbol{z^{id}}$, a face image $I_{face}$ and a non-face image $I_{nonface}$ as input, producing the positions $X$, color $C$, scale $S$, rotation $Q$ and opacity $A$ of the 3D Gaussians. Similar to the guiding geometry model, we initially maintain an overall mean point cloud, with the mean positions $\boldsymbol{X_{0}}$. However, we no longer generate the per-vertex feature $\Gamma$ through $\boldsymbol{f_{mean}}(x)$. Instead, we directly bind it to the Gaussian per-point feature as optimizable variables $\boldsymbol{\Gamma_{0}}$. This is possible since the number of Gaussian points is fixed at this stage. Then we need to transform the mean point cloud into a neutral-expression point cloud, conditioned by the id code $\boldsymbol{z^{id}}$. To achieve this, we utilize the same injection MLP $\boldsymbol{f_{inj}}(\cdot)$ and identity deformation MLP $\boldsymbol{f_{id}}(\cdot)$ defined in the guiding geometry model, which can generate feature vectors $H=\boldsymbol{f_{inj}}(\boldsymbol{z^{id}},\boldsymbol{\Gamma_{0}})$ that encode identity information for each point and predict the identity-related displacement $\delta X_{id}$ of each point. Then, we also need to predict the facial expression conditioned displacement $\delta X_{exp}$, the non-face displacement $\delta X_{nf}$, the resulting positions $X_{can}$ and the 32-channel color $C$ of each point, similar to the approach presented in the guiding geometry model. These can be described as:
$\delta X_{id}=$ $\boldsymbol{f_{id}}(H),$ (5)
$\delta X_{exp}=$ $\boldsymbol{f_{exp}}(H,\boldsymbol{E}_{exp}(I_{face})),$ (6)
$\delta X_{nf}=$ $\boldsymbol{f_{nf}}(H,\boldsymbol{E}_{nf}(I_{nonface})),$ (7)
$X_{can}=$ $\boldsymbol{X_{0}}+\delta X_{id}+\lambda_{exp}(\boldsymbol{X_{0}})\delta X_{exp}+$ (8)
$\lambda_{nf}(\boldsymbol{X_{0}})\delta X_{nf},$ (9)
$C=$ $\boldsymbol{f_{col}}(H).$ (10)